The smallest angle in a triangle is 1/3 of the largest angle. The third angle is 10 degree's more than the smallest angle. Find the measure of all three angles.
Answer by mananth(16946)

let largest angle be x
smallest angle = x/3
third angle = x/3 +10
..
x/3 +10 + x/3 + x =180
add -10 to both sides
and reorder
x+2x/3 = 180-10
5x/3 = 170
5x= 510
divide by 5
5x/5 = 510/5
x=102.
1/3 * 102 = 34 the second angle
third angle = x/3 + 10 = 34+10=44
the angles are 102, 34, 44